Good day,

Please Consider the reaction
CO(g)+NH3(g)⇌HCONH2(g), Kc=0.640
If a reaction vessel initially contains only CO and NH3 at concentrations of 1.00 M and 2.00 M, respectively, what will the concentration of HCONH2 be at equilibrium?

Thanks
1
Expert's answer
2020-08-09T08:42:47-0400

Consider the reaction

CO(g)+NH3(g)⇌HCONH2(g), Kc=0.640

If a reaction vessel initially contains only CO and NH3 at concentrations of 1.00 M and 2.00 M, respectively, what will the concentration of HCONH2 be at equilibrium?


Solution:

According to given reaction, we can make a table of concentrations for each compounds:

CO(g)+NH3(g)⇌HCONH2(g)

Unitial 1.00 2.00 0

React. x x x

Equilibrium 1.00-x 2.00-x x

The Kc has following equation:

Kc = [HCONH2]/([CO]*[NH3])

Using our values:

Kc = x/((1-x)*(2-x))

or:

x/((1-x)*(2-x)) = 0.640

x/(2-3x+x2) = 0.640

x = 0.640x2 - 1.92x + 1.28

0.640x2 - 2.92x + 1.28 = 0

where

x = 0.491

So the equilubrium concentrarion of HCONH2 equals 0.491 M


Answer: [HCONH2]eq = 0.491 M.